OpenStudy (anonymous):

Please help me! Given the equation 2 Square root of x minus 5= 2, solve for x and identify if it is an extraneous solution.

OpenStudy (anonymous):

\[This is the equation?:2\sqrt{x}-5=2\]

OpenStudy (anonymous):

no the square root is over the X and the 5

OpenStudy (anonymous):

@ifrimpanainte

OpenStudy (anonymous):

\[2\sqrt{x-5}=2|^{2}=>4(x-5)=4/:4=>x-5=1=>x=6\]

OpenStudy (anonymous):

thank you!